OpenStudy (anonymous):

A rate of 40 percent could be expressed as 2/5.

OpenStudy (anonymous):

yes because .4 = 4/10 = 2/5

jhonyy9 (jhonyy9):

yes but real is 40 percent can being write like 40/100 =4/10 =2/5